# Knowledge Graph Guided Evaluation of Abstention Techniques

### v1

@@ Abstract
To deploy language models safely, it is crucial that they abstain from responding to inappropriate requests. Several prior studies test the safety promises of models based on their effectiveness in blocking malicious requests. In this work, we focus on evaluating the underlying techniques that cause models to abstain. We create SELECT, a benchmark derived from a set of benign concepts (e.g., “rivers”) from a knowledge graph. The nature of SELECT enables us to isolate the effects of abstention techniques from other safety training procedures, as well as evaluate their generalization and specificity. Using SELECT, we benchmark different abstention techniques over six open-weight and closed-source models. We find that the examined techniques indeed cause models to abstain with over $80\%$ abstention rates. However, these techniques are not as effective for descendants of the target concepts, with refusal rates declining by $19\%$. We also characterize the generalization-vs-specificity trade-offs for different techniques. Overall, no single technique is invariably better than the others. Our findings call for a careful evaluation of different aspects of abstention, and hopefully inform practitioners of various trade-offs involved.
@@
Kinshuk Vasisht Indian Institute of Science Bengaluru, India kinshukv@iisc.ac.in                        Navreet Kaur thanks:  Work done as a Research Associate at Indian Institute of Science. University of Washington Seattle, USA kanavr@uw.edu                        Danish Pruthi Indian Institute of Science Bengaluru, India danishp@iisc.ac.in
@@ 1 Introduction
For several reasons, it is critical that language models abstain from responding to inappropriate user requests. These requests could include (malicious) attempts to assist users in illegal activity (Fang et al., 2024; Weidinger et al., 2021), generate offensive content (Deshpande et al., 2023; Gehman et al., 2020), or disseminate large-scale misinformation (Tamkin et al., 2021; Buchanan et al., 2021). To block such requests, a common approach is to perform some form of “safety” training before releasing language models (Bai et al., 2022). Safety training methods encompass supervised fine-tuning (SFT) (Zhou et al., 2023) and reinforcement learning (Ouyang et al., 2022; Rafailov et al., 2023) to align model outputs with human preferences.
@@
In our paper, we evaluate the effectiveness, generalization and specificity of abstention techniques. We introduce a benchmark, SELECT (Selective-abstention Evaluated by Leveraging an Extensive Concept Taxonomy), which derives benign concepts from a knowledge graph (KG)  (Suchanek et al., 2024). By focusing on benign concepts, we intend to isolate the effects of safety-training approaches performed for sensitive topics. Further, the connection with KGs allows us to exploit the hierarchical relations between entities to evaluate generalization and specificity.
@@
We evaluate abstention techniques including prompting (Zheng et al., 2024), activation steering (Lee et al., 2024) and fine-tuning (Bianchi et al., 2024; Rafailov et al., 2023) for both closed-source and open-weight models, including GPT-4o (OpenAI, 2024) and LLaMA-3.1 (Dubey et al., 2024). Through our evaluation, we find that abstention techniques are effective at introducing abstention for benign concepts and their compositions, with effectiveness as measured using abstention rates being over $80\%$ for most models, further reaching close to $100\%$ in some cases. However, the abstention effects do not necessarily extend to descendants, where we observe a drop of $19\%$ in abstention rates on average. Interestingly, different abstention techniques trade-off between generalization and specificity in varied ways. Most techniques perform similarly in terms of effectiveness corresponding to concepts at varying depths in the taxonomy. Somewhat surprisingly, specificity decreases for concepts that are farther from the root, suggesting that abstention for narrower concepts leads to over-refusal.
@@ 2 Related Work
When deploying large language models, it is crucial to abstain from inappropriate or malicious user requests (Weidinger et al., 2021; Fang et al., 2024). User requests—and corresponding model responses—may lead to bias, discrimination or toxicity (Deshpande et al., 2023; Gehman et al., 2020). Further, users may request information that is not captured within the parametric knowledge of the model (Feng et al., 2024), responses to which may be used to disseminate misinformation (Tamkin et al., 2021; Buchanan et al., 2021). Several recent efforts benchmark different facets of abstention in language models. For instance, SORRY-Bench (Xie et al., 2024) focuses on benchmarking the ability of models to abstain from user requests considered harmful and malicious. The study benchmarks various open and closed source models, finding closed-source models display tolerable refusal rates across different categories of harm. Further, UnknownBench (Liu et al., 2024) evaluates the ability of models to abstain in scenarios where models do not possess adequate knowledge to answer a question. The study notes that the refusal rates are far from perfect for GPT-4 models. Another benchmark, Priv-QA (Chen et al., 2023), evaluates abstention over protected groups whose information is considered confidential. Additionally, OR-Bench (Cui et al., 2024) studies the behavior of over-refusal prevalent in models aligned for safety, using a benchmark comprising toxic and seemingly-toxic but benign prompts. They conclude that models which effectively abstain from toxic prompts also abstain from the seemingly-toxic ones. A more comprehensive effort provides an extensive taxonomy of scenarios where abstention is desirable, and finds that open-source models are less safe compared to closed-source models (Wang et al., 2024). It is important to note that all the above studies focus on models rather than the underlying abstention techniques. Further, these benchmarks test models on issues that safety training attempts to address, thus it is difficult to isolate the effect of abstention techniques. By considering benign concepts, we circumvent this limitation.
@@ 3.1 Dataset Construction
To build SELECT, we identify five top-level entities in YAGO’s ontology as root concepts for our taxonomy: Products, People, Places, Organizations, and Creative Works. Most entities in YAGO are descendants of these concepts. For each concept, we recursively include descendants related by an ‘IS-A’ relation in YAGO. We repeat this process until we reach a leaf node, or a depth of six (whichever criteria is met first). For leaf nodes, such as “rivers”, we sample up to $5$ instances of the leaf concept. In order to study generalization and specificity, we altogether discard concepts that lie on short paths, wherein the distance between the root node and the leaf node is lower than $3$. This process results in a total of $394$ atomic concepts and instances.
@@
After collating the concepts, we use a language model (GPT-4o, OpenAI (2024)) to generate $50$ concept-related questions. We explicitly instruct the model to generate diverse questions about the concept that users are likely to ask. Along with our instruction, we also include five examples of user queries from WildChat-1M which is a repository of human-ChatGPT interactions (Zhao et al., 2024). We manually examine $180$ generated questions from $18$ randomly sampled concepts, and find them to be plausible and relevant to the concept in question. From the $50$ questions that we generate for every concept, we use $20$ for training (or few-shot examples), and reserve the remaining $30$ for evaluation. In case of compositions, we generate $30$ questions per concept and reserve $20$ for evaluation. In total, we obtain $11,820$ questions for evaluation for atomic concepts and $3,120$ for compositions—a sample of concepts, compositions and queries from SELECT is available in Table 1.
@@ 4.1 Abstention Techniques
Recent studies show that activation steering i.e. modifying model activations during inference can cause selective abstention (Turner et al., 2024; Zou et al., 2023a). We use conditional activation steering similar to CAST (Lee et al., 2024) to direct models towards abstaining from specific concepts. This technique requires access to the model’s weights to derive two vectors: concept and refusal. Concept vectors enable us to determine if a query is related to a given concept, and a refusal vector helps in steering the model to abstain from that concept. We derive the concept vector using questions from SELECT that are related and unrelated to a concept. Specifically, we select the first principal component of the difference in model activations for related and unrelated questions as the concept vector. We then classify new questions as concept-related if its cosine similarity with the concept vector lies above a threshold at specific layers.222 We use $20$ related and $20$ unrelated questions, and reserve $5$ from each for validation. Considering pairwise combinations, this leaves us with $225$ training and $25$ validation examples. The layers and thresholds for classifying concept-related questions are determined using the validation examples. Similarly, we obtain a refusal vector using the dataset from Zou et al. (2023a). We then steer the model to abstain from the target concept by adding a scaled refusal vector to selected layers, with the scaling factor and layers determined through trial and error (details in §C.2).
@@
We experiment with two fine-tuning approaches generally used for alignment: Supervised Fine-Tuning (SFT) (Zhou et al., 2023), and SFT followed by Direct Preference Optimization (DPO) (Rafailov et al., 2023). We sample $20$ questions about the target concept and $20$ for descendants, and generate $8$ abstention responses per question. To avoid over-refusal due to fine-tuning on abstained responses alone (Bianchi et al., 2024), we also include regular responses from outside the sub-tree, using $1,920$ instances for fine-tuning in total. Further, we construct abstention-compliance response pairs per question, with a preference for abstention on the target concept or its descendants, and a preference for compliance for the remainder, resulting in a total $960$ pairwise comparisons.333 We find a ratio of $1$:$5$ for abstention and compliance instances works best for SFT, and $1$:$2$ for SFT + DPO. Finally, we fine-tune the model for $10$ epochs using $5\times 10^{-5}$ learning rate, $16$ batch size, and the Adam optimizer (Kingma and Ba, 2017) with $0.01$ weight decay. Keeping all the hyperparameters same as SFT, we apply SFT + DPO for $5$ epochs using a smaller learning rate of $5\times 10^{-6}$. Further details are available in Appendix §C.3.
@@ 5.2 How well do abstention techniques perform for composition of concepts?
Furthermore, we find that abstention techniques over-refuse more often for compositions of concepts. Compared to atomic concepts, generalization for compositions increases by $9\%$ on average, while specificity decreases by $5$-$29$%. Notably, for fine-tuning methods, we observe that generalization exceeds abstention rates but the corresponding specificity rates are $15$-$29$% lower relative to atomic concepts, suggesting over-refusal. For compositions of concepts, activation steering has the best generalization-specificity trade-off, with a difference of $10\%$ compared to other techniques with larger gaps of $18$-$43$%. SFT demonstrates the worst trade-off across models, with specificity only at $47.7\%$ on average and a gap of $42.6\%$ between specificity and generalization, which is in stark contrast to the trends in the atomic concept setting where SFT offers the best balance. The large trade-off gap and the low specificity rates ($<70\%$ on average) suggest that current abstention techniques are not very effective for selective abstention from composition of concepts. The exact generalization and specificity rates for different models and abstention techniques over compositions are listed in Table 8 in the Appendix.
@@ 6 Conclusion
Limitations
We identify some important limitations of our work. We only consider a limited set of concepts as part of the benchmark, owing to decisions that prioritize a higher depth of the taxonomy that we curate. A lot of popular concepts and instances are, as a result, not a part of our taxonomy. The number of compositions we derive are limited further, as we rely on manually-curated templates and hand-crafted rules to compose concepts.
We employ a language model, notably GPT-4o, to generate questions related to concepts and compositions that are ultimately used for evaluation. While we attempt to validate whether these questions are plausible and related to the target concept they are derived from, it may not represent the distribution of actual user queries.
@@
Finally, we evaluate abstention only in a single-turn setting. In multi-turn conversations, user requests may include a mix of questions which must be selectively abstained and answered. Evaluations on recent benchmarks such as Fractured SORRY-Bench (Priyanshu and Vijay, 2024) demonstrates the vulnerabilities of safety-aligned models to inappropriate requests across multi-turn conversations. Our evaluations lack how abstention techniques behave when subject to adversarial perturbations (jailbreaks), or questions with a personalized context. We aim to target such evaluations in future.
@@
Our evaluations involve questions about benign and abstract concepts, rather than “sensitive” topics. The concepts are derived from YAGO, which exists in public domain (CC by 4.0)  [^4] https://creativecommons.org/licenses/by/4.0/, and the curation process is entirely automated—except the curation of templates for compositions. We plan to release all data and evaluation code to facilitate further research.

### v2

@@ Abstract
To deploy language models safely, it is crucial that they abstain from responding to inappropriate requests. Several prior studies test the safety promises of models based on their effectiveness in blocking malicious requests. In this work, we focus on evaluating the underlying techniques that cause models to abstain. We create SELECT, a benchmark derived from a set of benign concepts (e.g., “rivers”) from a knowledge graph. Focusing on benign concepts isolates the effect of safety training, and grounding these concepts in a knowledge graph allows us to study the generalization and specificity of abstention techniques. Using SELECT, we benchmark different abstention techniques over six open-weight and closed-source models. We find that the examined techniques indeed cause models to abstain with over $80\%$ abstention rates. However, these techniques are not as effective for descendants of the target concepts, where abstention rates drop by $19\%$. We also characterize the generalization-specificity tradeoffs for different techniques. Overall, no single technique is invariably better than others, and our findings inform practitioners of the various trade-offs involved. [^1] The benchmark and code to reproduce the evaluation with SELECT is available at https://github.com/kinshuk-h/SELECT
@@
Kinshuk Vasisht Indian Institute of Science Bengaluru, KA, India kinshukv@iisc.ac.in                        Navreet Kaur thanks:  Work done while at Indian Institute of Science. University of Washington Seattle, USA kanavr@uw.edu                        Danish Pruthi Indian Institute of Science Bengaluru, KA, India danishp@iisc.ac.in
@@ 1 Introduction
For several reasons, it is critical that language models abstain from responding to inappropriate user requests. These requests could include (malicious) attempts to assist users in illegal activity (Fang et al., 2024; Weidinger et al., 2021), generate offensive content (Deshpande et al., 2023; Gehman et al., 2020), or disseminate large-scale misinformation (Tamkin et al., 2021; Buchanan et al., 2021). To block such requests, a common approach is to perform some form of “safety” post-training before releasing language models (Bai et al., 2022). Safety post-training methods encompass supervised fine-tuning (SFT) (Zhou et al., 2023) and reinforcement learning (Ouyang et al., 2022; Rafailov et al., 2023) to align model outputs with human preferences.
@@
In this work, we introduce SELECT (Selective-abstention Evaluated by Leveraging an Extensive Concept Taxonomy), a benchmark to evaluate the effectiveness, generalization and specificity of abstention techniques. SELECT derives benign concepts from a knowledge graph (KG)  (Suchanek et al., 2024). By focusing on benign concepts, we intend to isolate the effects of safety-training approaches performed for sensitive topics. Further, the connection with KGs allows us to exploit the hierarchical relations between entities to evaluate generalization and specificity.
@@
Using SELECT, we evaluate abstention techniques including prompting (Zheng et al., 2024), activation steering (Lee et al., 2024) and fine-tuning (Bianchi et al., 2024; Rafailov et al., 2023) for both closed-source and open-weight models, including GPT-4o (OpenAI, 2024) and LLaMA-3.1 (Dubey et al., 2024). Through our evaluation, we find that abstention techniques are effective for abstention from benign concepts and their compositions, measured using abstention rates being over $80\%$ for most techniques and further reaching close to $100\%$ in some cases. However, the abstention effects do not necessarily extend to descendants, where we observe abstention rates drop by $19\%$ on average. Interestingly, different techniques trade-off between generalization and specificity in varied ways. Most techniques perform similarly in terms of effectiveness corresponding to concepts at varying depths in the taxonomy. Somewhat surprisingly, specificity decreases for concepts that are farther from the root, suggesting that abstention for narrower concepts leads to over-refusal.
@@ 2 Related Work
When deploying large language models, it is crucial to abstain from inappropriate or malicious user requests (Weidinger et al., 2021; Fang et al., 2024). User requests—and corresponding model responses—may lead to bias, discrimination or toxicity (Deshpande et al., 2023; Gehman et al., 2020). Further, users may request information that is not captured within the parametric knowledge of the model (Feng et al., 2024), responses to which may be used to disseminate misinformation (Tamkin et al., 2021; Buchanan et al., 2021). Several recent efforts benchmark different facets of abstention in language models. For instance, SORRY-Bench (Xie et al., 2024) focuses on benchmarking the ability of models to abstain from user requests considered harmful and malicious. The study benchmarks various open and closed source models, finding closed-source models display tolerable refusal rates across different categories of harm. Further, UnknownBench (Liu et al., 2024) evaluates the ability of models to abstain in scenarios where models do not possess adequate knowledge to answer a question. The study notes that the refusal rates are far from perfect for GPT-4 models. Another benchmark, Priv-QA (Chen et al., 2023), evaluates abstention over protected groups whose information is considered confidential. Additionally, XS-Test (Röttger et al., 2024) and OR-Bench (Cui et al., 2024) study the behavior of over-refusal prevalent in models aligned for safety, using a benchmark comprising toxic and seemingly-toxic but benign prompts. Their findings conclude that models which effectively abstain from toxic prompts also abstain from the seemingly-toxic ones. Other benchmarks such as Do-Not-Answer (Wang et al., 2024) and CoCoNot (Brahman et al., 2024) conduct a more comprehensive evaluation based on an extensive taxonomy of scenarios where abstention is desirable. Using Do-Not-Answer, the authors find that open-source models are less safe compared to closed-source models (Wang et al., 2024). The taxonomy of CoCoNot further incorporates incomplete, unsupported and indeterminate requests to evaluate a broader notion of non-compliance. Their analysis finds that while compliance for unsafe requests is low, other categories show higher degrees of compliance (Brahman et al., 2024). It is important to note that all the above studies focus on models rather than the underlying abstention techniques. Further, these benchmarks test models on issues that safety training attempts to address, thus it is difficult to isolate the effect of abstention techniques. By considering benign concepts, we circumvent this limitation.
@@ 3.1 Dataset Construction
To build SELECT, we identify five top-level entities in YAGO’s ontology as root concepts for our taxonomy: Products, People, Places, Organizations, and Creative Works. Most entities in YAGO are descendants of these concepts. For each concept, we recursively include descendants related by an ‘IS-A’ relation in YAGO. We repeat this process until we reach a leaf node, or a depth of six (whichever criteria is met first). For leaf nodes, such as “rivers”, we randomly sample up to $K$ instances of the leaf concept. In our setup, we choose $K=5$ due to higher computational overhead associated with processing a larger number of leaf nodes. In order to study generalization and specificity, we altogether discard concepts that lie on short paths, wherein the distance between the root node and the leaf node is lower than $L$. We choose $L=3$, as smaller values limit the depth of the taxonomy, while larger values may prune out popular concepts that are at shallow depths from the root. This process results in a total of $394$ atomic concepts and instances.
@@
Characteristic Atomic Concepts Composition of Concepts
Number of concepts 394 156
Maximum depth 6 5
Number of leaves 290 98
Children per non-leaf (avg.) 3.7 2.4
Ancestors per node (avg.) 3.2 1.7
Number of questions $11,820$ $3,120$
Lexical diversity (TTR) $0.61$ $0.57$
Table 2: Characteristics of the data comprising SELECT. We measure the lexical diversity of questions using Type-Token Ratio (TTR) – the ratio of unique words to words (Johnson, 1944).
After collating the concepts, we use a language model (GPT-4o, OpenAI (2024)) to generate $50$ concept-related questions. We explicitly instruct the model to generate diverse questions about the concept that users are likely to ask. For concepts such as ‘Washington’ which may refer to either a place or a person, we disambiguate the context by specifying the list of ancestors alongside the concept in the prompt (e.g., ‘Washington in the context of Places, Administrative Areas, States’). This is also followed when employing the abstention technique. Along with our instruction, we also include five examples of user queries from WildChat-1M which is a repository of human-ChatGPT interactions (Zhao et al., 2024). We manually examine $180$ generated questions from $18$ randomly sampled concepts, and find them to be plausible and relevant to the concept in question. From the $50$ questions that we generate for every concept, we use $20$ for training (or few-shot examples), and reserve the remaining $30$ for evaluation. In case of compositions, we generate $30$ questions per concept and reserve $20$ for evaluation. In total, we obtain $11,820$ questions for evaluation for atomic concepts and $3,120$ for compositions—a sample of concepts, compositions and queries from SELECT is available in Table 1. Other characteristics of the data are summarized in Table 2.
@@ 4.1 Abstention Techniques
Recent studies show that activation steering i.e. modifying model activations during inference can cause selective abstention (Turner et al., 2024; Zou et al., 2023a). We use conditional activation steering similar to CAST (Lee et al., 2024) to direct models towards abstaining from specific concepts. This technique requires access to the model’s weights to derive two vectors: concept and refusal. Concept vectors enable us to determine if a query is related to a given concept, and a refusal vector helps in steering the model to abstain from that concept. We derive the concept vector using questions from SELECT that are related and unrelated to a concept. Specifically, we select the first principal component of the difference in model activations for related and unrelated questions as the concept vector. We then classify new questions as concept-related if its cosine similarity with the concept vector lies above a threshold at specific layers.333 We use $20$ related and $20$ unrelated questions, and reserve $5$ from each for validation. Considering pairwise combinations, this leaves us with $225$ training and $25$ validation examples. The layers and thresholds for classifying concept-related questions are determined using the validation examples. Similarly, we obtain a refusal vector using the dataset from Zou et al. (2023a). We then steer the model to abstain from the target concept by adding a scaled refusal vector to selected layers, with the scaling factor and layers determined through trial and error (details in §C.2).
@@
We experiment with two fine-tuning approaches generally used for alignment: Supervised Fine-Tuning (SFT) (Zhou et al., 2023), and SFT followed by Direct Preference Optimization (DPO) (Rafailov et al., 2023). We sample $20$ questions about the target concept and $20$ for descendants, and generate $8$ abstention responses per question. To avoid over-refusal due to fine-tuning on abstained responses alone (Bianchi et al., 2024), we also include regular responses from outside the sub-tree, using $1,920$ instances for fine-tuning in total. Further, we construct abstention-compliance response pairs per question, with a preference for abstention on the target concept or its descendants, and a preference for compliance for the remainder, resulting in a total $960$ pairwise comparisons.444 We find a ratio of $1$:$5$ for abstention and compliance instances works best for SFT, and $1$:$2$ for SFT + DPO. Finally, we fine-tune the model for $10$ epochs using $5\times 10^{-5}$ learning rate, $16$ batch size, and the Adam optimizer (Kingma and Ba, 2017) with $0.01$ weight decay. Keeping all the hyperparameters same as SFT, we apply SFT + DPO for $5$ epochs using a smaller learning rate of $5\times 10^{-6}$. Further details are available in Appendix §C.3.
@@ 5.2 How well do abstention techniques perform for composition of concepts?
Furthermore, we find that abstention techniques over-refuse more often for compositions of concepts. Compared to atomic concepts, generalization for compositions increases by $9\%$ on average, while specificity decreases by $5$-$29$%. Notably, for fine-tuning methods, we observe that generalization exceeds abstention rates but the corresponding specificity rates are $15$-$29$% lower relative to atomic concepts, suggesting over-refusal. For compositions of concepts, activation steering has the best generalization-specificity trade-off, with a difference of $10\%$ compared to other techniques with larger gaps of $18$-$43$%. SFT demonstrates the worst trade-off across models, with specificity only at $47.7\%$ on average and a gap of $42.6\%$ between specificity and generalization, which is in stark contrast to the trends in the atomic concept setting where SFT offers the best balance. The large trade-off gap and the low specificity rates ($<70\%$ on average) suggest that current abstention techniques are not very effective for selective abstention from composition of concepts. The exact generalization and specificity rates are listed in Table 9.
5.3 Discussion
Our findings shed light on the performance and trade-offs of different abstention techniques. We find that no technique is invariably better than the others. Further, the choice of an abstention technique for a given scenario is motivated by several factors. These factors include the granularity of the concept, desired levels of specificity and generalization, and the compute and data constraints.
We believe our findings have important implications for abstaining from novel unsafe concepts. Situations that require abstention to generalize (e.g., new forms of malicious use, toxicity, etc.) can benefit from fine-tuning methods for abstention, given the fact that their generalization capabilities are consistent across different granularities of concepts. On the other hand, when abstaining from fine-grained concepts (such as misinformation about a topic or discrimination towards a specific social group), techniques with higher specificity such as activation steering are recommended.
@@ 6 Conclusion
Limitations & Future Work
We identify some important limitations of our work. We employ a language model, notably GPT-4o, to generate questions related to concepts and compositions that are ultimately used for evaluation. We find these questions to be lexically diverse, and a subset of these questions to be plausible and related to the target concept. However, these questions may not be representative of actual user queries. Further, as we later evaluate GPT-4o with the same questions, the results for the model may be biased.
@@
Safety training is also known to be suspectible to jailbreaks (Wei et al., 2023). Our preliminary analysis shows that abstentions techniques we study can also be fragile (§E.5). Future work can explore the adversarial robustness of abstention techniques in more depth, and ways to improve them.
Lastly, real-world user interactions are generally multi-turn and multilingual. Recent studies highlight the fragility of safety training in multi-turn and multilingual conversations (Priyanshu and Vijay, 2024; Poppi et al., 2024). Evaluating the effectiveness of abstention techniques in this challenging setting, where models may be required to selectively abstain and comply additionally in multiple languages, is an exciting future direction.
@@
Our evaluations involve questions about benign and abstract concepts, rather than “sensitive” topics. The concepts are derived from YAGO, which exists in public domain (CC by 4.0)  [^5] https://creativecommons.org/licenses/by/4.0/, and the curation process is entirely automated—except the curation of templates for compositions. We release all data and evaluation code to facilitate further research.
@@
Acknowledgements
We thank the anonymous reviewers for their feedback. We also thank Mitesh Khapra, Miroojin Bakshi, Badrinath Singhal, Priyam Dey and A.S. Anudeep for their suggestions and insights. DP is grateful to Adobe Inc., Google Research, and the Kotak IISc AI-ML Centre (KIAC) for generously supporting his group’s research.